A mechanic sells a brand of automobile tire that has a life expectancy that is normally​ distributed, with a mean life of 29 comma 000 miles and a standard deviation of 2800 miles. He wants to give a guarantee for free replacement of tires that​ don't wear well. How should he word his guarantee if he is willing to replace approximately​ 10% of the​ tires?

Tires that wear out by [ ] miles will be replaced free of charge.

​(Round to the nearest mile as​ needed.)
Answer #1

Given that,

mean = = 29000

standard deviation = = 2800

Using standard normal table,

P(Z < z) =10 %

= P(Z < z) = 0.10  

= P(Z <-1.28 ) = 0.10

z = -1.28 Using standard normal z table,

Using z-score formula  

x= z * +

x= -1.28 *2800+29000

x= 25416